Solve:


x².d²y/dx² -x.dy/dx +y=ln x

1
Expert's answer
2022-02-22T10:47:11-0500

Let us use the transformation "x=e^t." Then 

"y'_x=y'_te^{-t},\ y''_{x^2}=(y''_{t^2}-y'_t)e^{-2t}."


We get the equation for the function "y(t)"

"e^{2t}(y''-y')e^{-2t} -e^ty'e^{-t} +y=t,"

 which is equivalent to 

"y''-2y'+y=t"


Corresponding homogeneous differential equation


"y''-2y'+y=0"

Characteristic (auxiliary) equation


"r^2-2r+1=0"

"r_1=r_2=1"

The general solution of the homogeneous differential equation is


"y_h=c_1e^t+c_2 te^t"


Find the particular solution of the non homogeneous differential equation


"y_p=At+B"

"y_p '=A"

"y''_p=0"

Substitute


"0-2A+At+B=t"

"A=1, B=2"

The general solution of the non homogeneous differential equation is


"y=c_1e^t+c_2 te^t+t+2"

"y(x)=c_1x+c_2x\ln x+\ln x+2"